Nichols Inc. manufactures remote controls. Currently the company uses a

plantminus−wide

rate for allocating manufacturing overhead. The plant manager is considering

switchingminus−over

to ABC costing system and has asked the accounting department to identify the primary production activities and their cost drivers which are as​ follows:

Activities

Cost driver

Allocation Rate

Material handling

Number of parts

​$2 per part

Assembly

Labor hours

​$20 per hour

Inspection

Time at inspection station

​$3 per minute

The current traditional cost method allocates overhead based on direct manufacturing labor hours using a rate of​ $200 per labor hour.

What are the indirect manufacturing costs per remote control assuming an

activityminus−basedminus−costing

method is used and a batch of 50 remote controls are​ produced? The batch requires 100​ parts, 6 direct manufacturing labor​ hours, and 2.5 minutes of inspection time.

A.

​$4.00 per remote control

B.

​$327.50 per remote control

C.

​$6.55 per remote control

D.

​$24.00 per remote control
Answer #1
Material handling 200 =100*2
Assembly 120 =6*20
Inspection 7.5 =2.5*3
Total indirect manufacturing costs 327.5
Divide by Total remote controls 50
Indirect manufacturing costs per remote control 6.55
Option C $6.55 per remote control is correct